Eva invested a certain amount of money at 8% interest and $1600 more than that amount at 12%. Her total yearly interest was $845. How much did she invest at each rate?
Answer by stanbon(75887)

Eva invested a certain amount of money at 8% interest and $1600 more than that amount at 12%. Her total yearly interest was $845. How much did she invest at each rate?
--------
Interest Equation:
0.08x + 0.12(x+1600) = 845
--------
8x + 12x + 12*1600 = 84500
20x = 65300
x = 3265 dollars (amt. invested at 8%)
x+1600 = 4865 dollars (amt. invested at 12%)
=================
Cheers,
Stan H.
=================